# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ Abstract
In this paper, we extend two classes of functions classically involved in asymptotic stability analyses for studying a maximization problem on the reachable values of a discrete-time dynamical system. This particular maximization problem is called a peaks computation problem. The problem is to find a couple composed of an initial state and a time which maximizes a given function over states. The paper focuses on the time component of the optimal solution which is an integer as the time is discrete. We develop a method to provide an upper bound of this integer from a formula which requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. A first result proves that the formula provides, in theory, the optimal integer. However, in practice, the computation cannot be so precise. Then, we develop two alternative methods. The first is based on discontinuous and non strictly increasing/decreasing $\mathcal{KL}$-like functions named $\mathcal{KL}_{\rm gen}$ functions. We prove that the existence of a $\mathcal{KL}_{\rm gen}$ upper bound is equivalent to the existence of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of the strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function needs an extension of the famous Sontag’s lemma. Finally, we construct a new type of Lyapunov functions well designed for our peaks computation problem. These functions are called Opt-Lyapunov functions. Opt-Lyapunov functions are well designed since we establish an equivalence theorem between the existence of an Opt-Lyapunov function and of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of a Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is insipred by the Yoshizawa construction of Lyapunov functions.
Keywords: Discrete-time, Peaks Computation Problem, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
@@ 1 Introduction
In this paper, we study the problem of peaks computation for discrete-time dynamical systems in finite dimensional state spaces. A peaks computation problem can be viewed as a particular maximization problem for which the constraints set is the reachable values set of a discrete-time dynamical system. The objective function of this optimization problem is a function over the states. This particular maximization problem can model various situations in engineering, economics, physics, biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species even if the system asymptotically oscillates between several equilibra. This problem boils down to maximize separetely each coordinate of the state variable over the system modeling the population dynamics.
@@
Besides its importance to analyze critical concrete situations, the literature about peaks analysis problem in the community of dynamical systems analysis is pretty thin. The peaks computation problem appears in a discrete-time setting for linear systems and where the objective function is the Euclidean norm [APS18]. The problem also appears for the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj23]. A slighty different version of the problem is also encountered in [AG24]. The difference comes from the fact that the state variable is, moreover, constrained to stay in a some given set. The problem is studied in a continuous-time setting and in, algebraically, a polynomial environment in [MHS20]. Note that in [MHS20], the authors address the problem of verifying safety properties (see e.g. [BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peaks computation problem. Safety verification problems aim to prove that the trajectories of the system are contained in some safe set. A safety verification problem can be translated into a peaks computation problem when the safe set is a sublevel of a sufficiently regular function. In this case, the objective function of the peaks computation problem is the function associated with the sublevel set. The safety analysis from the peaks computation point-of-view is also proposed for switched system [Adj17] and for program analysis in [AGM15]. For safety analysis of dynamical systems, one can think that a reachability analysis (see e.g. [RKML06] and reference therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties on the dynamical system. However, maximality in our context is a synonym for criticality. The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition where the system could enter into.
@@
In [MHS20] the authors are interested in computing an upper bound of the peaks optimal value for the continuous case. The techniques used is based on Sums-Of-Squares (SOS) and Linear Matrix Inequalities (LMI) formulations (the interested reader can consult [Las15] for those subjects). The same goal motivates the authors of [APS18] and, as they limit their study to linear system and norm objective functions, they directly use semidefinite programming (see [BTN01] for further details on the subject). The reference [AG24] cannot be fairly compared due to the presence of the additional constraint. Moreover, again, the studied systems are linear or piecewise linear and the objective functions are linear. The goal of the current paper is to provide the theoretical bases to solve exactly a peaks computation problem in a general way. The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj23]
@@
In this paper, we propose to study peaks computation problems from classical tools of dynamical systems stability theory (e.g. [Ela05, Chap. 4] for discrete-time). This approach comes from a simple observation: the existence of peaks depends on the asymptotic behaviour of the system. The importance of stability analysis is time-tested and a stability analysis is a prerequisite to the implementation of numerical methods. Whereas general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g. [ATP10] or [NT04]) i.e. the norms of the states are globally bounded from above by a $\mathcal{KL}$ function. However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of state has to be done for all integers which can be difficult. On the other hand, Lyapunov functions are functional certificates for the system stability. Some results guarantee the existence of Lyapunov functions: a system is globally asymptotic stable if and only if there exists a Lyapunov function (see  e.g. [Kha02]). Being a Lyapunov function is a timeless property and only depends on dynamics. In few advantageous situations, Lyapunov functions can even be computed using numerical and optimization solvers (see the survey [GH15]).
@@
Formally, the data of a peaks computation problem is decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states. The dynamical system consists of a set of initial conditions $X^{\mathrm{i}n}$ and the dynamics of the system i.e. a function $T$ which maps a state to a new state. The associated peaks computation problem aims in maximizing $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{i}n}$ and $k$ is a natural integer. The optimal solution of a peaks computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer. Note that when an integer $k$ is fixed, maximizing $\varphi(T^{k}(x))$ over $x\in X^{\mathrm{in}}$ is a standard maximization problem. The main difficulty of peaks computation problem is to compute an optimal $k_{\rm opt}$ i.e. an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest. Our approach is to find an upper bound $K$ of the (greatest) integer maximizer. We impose to the upper bound $K$ to satisfy the following condition : for all $n$ greater than $K$, the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is smaller than the supremum of $\varphi(T^{K}(x))$. Once this integer $K$ computed, to solve the peaks computation problem can be done by solving a finite number of maximization problems. We can extract the optimal solution $(x_{\rm opt},k_{\rm opt})$ from one of those maximization problems. Thus we reduce the peaks computation problem to find the smallest possible integer $K$ satisfying the previous condition. We, then, provide a general formula to compute such integer. This formula relies on well-chosen pairs of a strictly increasing and continuous function on $[0,1]$, $h$, and a scalar $\beta\in(0,1)$. In this case, well-chosen means that the pair $(h,\beta)$ forms a sequence $(h(\beta^{k}))_{k}$ such that $(h(\beta^{k}))_{k}$ is greater than the sequence of the suprema of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. A well-chosen pair $(h,\beta)$ is said to be useful for our problem if $h(0)$ is strictly smaller than a supremum of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. We prove that there exists an optimal useful pair, in the sense of that the pair permits to find the greatest optimal integer. In practice, this pair is not computable. Therefore, we propose two other methods to construct useful pairs. First, we extend the class of $\mathcal{KL}$ functions to the class of functions which are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states and decreasing on time. We call this class of functions $\mathcal{KL}_{\rm gen}$. As for classical $\mathcal{KL}$ functions, we use $\mathcal{KL}_{\rm gen}$ functions as upper bounds on $\varphi(T^{k}(x))$ (for the classical case with $\mathcal{K}\mathcal{L}$, $\varphi$ is a norm). We establish an equivalence theorem where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$. The construction of useful pairs $(h,\beta)$ from $\mathcal{KL}_{\rm gen}$ upper bound functions relies on an extension of Sontag’s result [Son98, Prop. 7] for $\mathcal{KL}_{\rm gen}$. Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem namely the Opt-Lyapunov functions. The word adapted means the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$. Once again, the proofs of each implication are constructive. The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by Yoshizawa construction of Lyapunov functions [Yos66]. The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and to take infinity values. Those two new possibities make irrelevant the use of classical comparison functions (see e.g. [Kel14] for a comprehensive survey). In our context, when we have access to classical Lyapunov functions, we can link continuous $\varphi$ with classical Lyapunov functions through comparison functions. Without this link, we need to develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function. In this paper, we introduce the notion of certificate of compatibility to overcome the loss of the link.
In summary, the main contributions of the paper are the following:
  • we propose a novel approach of the peaks computation problem by bounding the number of static optimization problem to solve to obtain the optimal value;
  • we obtain a general upper bound method for bounded from above sequence based on homeomorphic deformations of geometric sequences;
@@
  • we generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence ;
@@
The paper is organized as follows. Section 2 presents the peaks computation problem addressed in this paper and the main difficulty related to the search of an integer as an optimal solution of the peaks computations problem. Section 3 recalls some basic results about the supremum of sequences and how it is related to their asymptotic behaviour. In Section 3, we also find the fundamental result about the existence of a useful pair of a strictly increasing and continuous function on $[0,1]$ and a scalar in $(0,1)$. This pair furnishes, in theory, the exact integer optimal solution. However, this pair is difficult to compute in practice. Then, in Section 4, after developing a simple result using the classical $\mathcal{K}\mathcal{L}$ class, we introduce the new class $\mathcal{KL}_{\rm gen}$. We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds of $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair. In Section 5, we present the concept of Opt-Lyapunov functions well designed for our maximization problem. The classical comparison functions associated to classical Lyapunov functions are replaced by a new notion called compatibility. We also prove that the existence of a compatible Opt-Lyapunov function is equivalent to the existence of a useful pair. In Section 6, we summarize the results obtained highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$ and convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class and Opt-Lyapunov functions.
@@ 2 Problem Statement And Motivated Example
2.1 Problem Statement
@@
Let $X^{\mathrm{in}}$ be a subset of $\mathbb{R}^{d}$ and $T$ a nonzero self-map on $\mathbb{R}^{d}$. Let us consider the associated discrete-time dynamical system defined as follows:
$x_{0}\in X^{\mathrm{in}},\ x_{k+1}=T(x_{k})\enspace.$
Let us consider $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$. We are interested in the reachable values that maximizes $\varphi$. More precisily, writing $T^{k}$ for the $k$-fold composition of $T$ or the identity when $k=0$, we are interested in solving the following maximization problem called a peaks computation problem:
Problem 1 (Peaks Computation Problem).
Maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
@@
Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ by $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure.
To solve Problem 1 means that we have to compute the optimal value $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$ and, if there exists, to characterize and to compute a couple of a vector $x\in X^{\mathrm{in}}$ and an integer $k$ which maximizes the value $\varphi(T^{k}(x))$. We can solve Problem 1 sequentially by solving a sufficient number of static maximization problems :
Problem Pk (Static Problems).
Maximize $\varphi(T^{k}(x))$ under the constraint $x\in X^{\mathrm{in}}$.
Some hypotheses can be made to ensure that Problem Pk has a finite optimal values and an optimal solution. For example, if $X^{\mathrm{in}}$ is compact and the function $\varphi\circ T^{k}$ is upper semicontinuous, Problem Pk has an optimal solution. About computations, some special instances of Problem Pk can be solved numerically more or less efficiently. For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, the problem is said to be convex and numerous methods exist (see e.g. [BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ a basic semialgebraic set (see e.g. [AL11, Las15]). The main issue remains to guess the interesting $k$, the integer for which the optimal value of Problem Pk is the greatest. A direct computation of such $k$ seems to be very difficult. Hence, the method we use in this paper is to solve instances of Problem Pk for $k=0,\ldots,K$ where $K$ has to determined. The integer $K$ has the particularity of ensuring that the optimal value of Problem 1 corresponds to the optimal value of some Problem Pk where $k$ is smaller than $K$.
In this paper, we will adopt the following notations:
$\nu:=(\nu_{k})_{k\in\mathbb{N}}\text{ where }\forall\,k\in\mathbb{N},\ \nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\text{ and }\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}$ (1)
Sometimes we will also use the terminlogy the supremum of $\nu$ to speak about $\nu_{\rm opt}$. The sequence $\nu$ corresponds to the sequence of optimal values of the maximization problems depicted at Problem Pk and $\nu_{\rm opt}$ is the optimal value of Problem 1. Following those notations, the main problem of the paper can be formulated mathematically as follows:
Problem 2 (Safe Number of Problem $P_{k}$ to Solve).
Find $K\in\mathbb{N}$ such that $\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}$.
In this paper, we provide simple characterizations for which a solution $K$ of Problem 2 exists. Then we construct a formula based on pairs composed of a strictly increasing continuous functions on $[0,1]$ and a scalar in $(0,1)$. Finally, we propose constructions of those strictly increasing continuous functions on [0,1] based on extensions of classical tools on stability theory for discrete-time dynamical systems such as $\mathcal{K}\mathcal{L}$ class of functions and Lyapunov functions.
The developments made in the paper require real-valued sequences; hence the only assumption made in this paper is the following :

### v4

# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peak Analysis

@@ Abstract
In this paper, we extend two classes of functions that are classically involved in asymptotic stability analyses.The purpose of this extension is to study a maximization problem on the reachable values of a discrete-time dynamical system. This particular maximization problem is called a peak computation problem. The problem consists in finding a couple composed of an initial state and a time that maximizes a given function over states. The paper focuses on the time component of the optimal solution, which is an integer as the time is discrete. We apply a method developed in previous papers to compute an upper bound of the greatest index maximizer of a real sequence. This previous method is based on a formula that requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence that provides an upper bound of the analyzed sequence. This pair is proven to exist. However, in practice, the computation cannot be done in a general setting. In this paper, we developed two alternative methods. The first is based on discontinuous and non-strictly increasing/decreasing $\mathcal{KL}$-like functions called $\mathcal{KL}_{\rm gen}$ functions. We prove that the existence of a $\mathcal{KL}_{\rm gen}$ upper bound is equivalent to the existence of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of a strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function requires an extension of the famous Sontag’s lemma. Finally, we construct a new type of Lyapunov function that is well designed for our peak computation problem. These functions are called Opt-Lyapunov functions. Opt-Lyapunov functions are suitable as we establish an equivalence theorem between the existence of an Opt-Lyapunov function and that of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of an Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is inspired by the Yoshizawa construction of Lyapunov functions.
Keywords: Discrete-time, Peak Computation Problems, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
@@ 1 Introduction
In this paper, we study the problem of peak computation for discrete-time dynamical systems in finite-dimensional state spaces. A peak computation problem can be viewed as a particular maximization problem for which the constraint set is the reachable value set of a discrete-time dynamical system. The objective function of this optimization problem is a function of the states. This particular maximization problem can be used to model various situations in engineering, economics, physics, and biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species, even if the system asymptotically oscillates between several equilibria. This problem boils down to maximizing each coordinate of the state variable separately over the system modeling the population dynamics.
@@
Besides its importance in analyzing critical concrete situations, the literature on peak analysis problems in the community of dynamical systems analysis is quite limited. The peak computation problem appears in a discrete-time setting for linear systems, where the objective function is the Euclidean norm [APS18]. This problem also appears in the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj25c]. A slightly different version of the problem was encountered in [AG24]. The difference arises from the fact that the state variable is constrained to remain in a given set. The problem was studied in a continuous-time setting and algebraically in a polynomial environment in [MHS20]. In  [MHS20], the authors address the problem of verifying safety properties (see e.g.,  [BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peak computation problem. Safety verification problems aim to prove that the trajectories of a system are contained in a safe set. A safety verification problem can be translated into a peak computation problem when the safe set is a sublevel of sufficiently regular function. In this case, the objective function of the peak computation problem is associated with the sublevel set. A safety analysis from the peak computation point of view was also proposed for switched systems  [Adj17] and for program analysis in [AGM15]. For the safety analysis of dynamical systems, one can think that a reachability analysis (see e.g.,  [RKML06] and references therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties of the dynamical system. However, maximality in our context is synonymous with criticality. The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition in which the system can enter.
@@
In [MHS20] the authors are interested in computing an upper bound of the peak optimal value for the continuous case. The techniques used are based on sum-of-squares (SOS) and linear matrix inequality (LMI) formulations (the interested reader can consult [Las15] for those subjects). The same goal motivated the authors of [APS18]. They limited their study to linear systems and norm objective functions and directly used semidefinite programming (see [BTN01] for further details on the subject). Reference  [AG24] cannot be fairly compared because of the additional constraint. Moreover, the studied systems are linear or piecewise linear, and the objective functions are linear. The goal of the current paper is to provide the theoretical bases to solve exactly a peak computation problem in a general way. The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj25c, Adj25a] and a specialization of [Adj25b] to specific sequences of optimal values arising in peak analysis computations.
@@
In this paper, we propose to study peak computation problems from extensions of classical tools of dynamical systems stability theory (for example,  [Ela05, Chap. 4] for discrete time). This approach is based on a simple observation: the existence of peaks depends on the asymptotic behavior of the system. The importance of stability analysis is time-tested, and stability analysis is a prerequisite for the implementation of numerical methods. Although general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g.,  [ATP10] or [NT04]), that is, the norms of the states are globally bounded from above by a $\mathcal{KL}$ function. However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of the state must be done for all integers, which can be difficult. In contrast, Lyapunov functions are functional certificates for system stability. Some results guarantee the existence of Lyapunov functions: a system is globally asymptotically stable if and only if there exists a Lyapunov function (see, e.g., [Kha02]). Being a Lyapunov function is a timeless property that depends only on the dynamics. In a few advantageous situations, Lyapunov functions can be computed using numerical and optimization solvers (see the survey [GH15]).
@@
Formally, the data of a peak computation problem are decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states. The dynamical system consists of a set of initial conditions $X^{\mathrm{in}}$ and the dynamics of the system, that is, a function $T$ that maps a state to a new state. The associated peak computation problem aims to maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k$ is a natural integer. An optimal solution of a peak computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer. Note that when an integer $k$ is fixed, maximizing $\varphi(T^{k}(x))$ over $x\in X^{\mathrm{in}}$ is a standard maximization problem. The main difficulty of the peak computation problem is to compute an optimal $k_{\rm opt}$, that is, an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest. The optimal value of a peak computation problem can be interpreted as the sequence of the optimal value of standard maximization problems where the integer $k$ is fixed. From this interpretation, the integer element $k_{\rm opt}$ of an optimal solution becomes an index for which the associated term maximizes the sequence. Therefore, we can apply the method developed in [Adj25b]. It consists in computing an upper bound of the (greatest) index maximizer of a real sequence that is bounded from above. The computation of this upper bound is based on a formula involving a pair $(h,\beta)$ where $h$ is a strictly increasing continuous function on $[0,1]$ and $\beta$ is a scalar belonging to $(0,1)$. The pair $(h,\beta)$ is also constrained to form a sequence $(h(\beta^{k}))_{k\in\mathbb{N}}$ greater than the sequence of the suprema of $\varphi\circ T^{k}$ over $X^{\mathrm{in}}$. To be applicable, the pair $(h,\beta)$ has to be useful, that is, $h(0)$ must be strictly smaller than, at least, a supremum of $\varphi\circ T^{k}$ over $X^{\mathrm{in}}$. In [Adj25b], we proved that such a pair $(h,\beta)$ exists and can be even chosen to be optimal, meaning that the pair permits finding the greatest optimal integer. In practice, this pair is not computable. Therefore, we propose two additional methods for constructing useful pairs. First, we extend the class of $\mathcal{KL}$ functions to the class of functions that are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states, and decreasing on time. We call this class of functions $\mathcal{KL}_{\rm gen}$. As for classical $\mathcal{KL}$ functions, we use $\mathcal{KL}_{\rm gen}$ functions as upper bounds on $\varphi(T^{k}(x))$ (for the classical case with $\mathcal{K}\mathcal{L}$, $\varphi$ is a norm). We establish an equivalence theorem, where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function, and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$. The construction of useful pairs $(h,\beta)$ from $\mathcal{KL}_{\rm gen}$ upper bound functions relies on an extension of Sontag’s result [Son98, Prop. 7] for $\mathcal{KL}_{\rm gen}$. Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem, namely, the Opt-Lyapunov functions. The word adapted means that the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$. Once again, the proofs of all implications are constructive. The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by the Yoshizawa construction of Lyapunov functions [Yos66]. The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and take infinite values. These two new possibilities make the use of classical comparison functions irrelevant (see e.g.,  [Kel14] for a comprehensive survey). In our context, when we have access to classical Lyapunov functions, we can link, when $\varphi$ is continuous, $\varphi$ with classical Lyapunov functions using comparison functions. Without this link, we must develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function. In this paper, we introduce the notion of a certificate of compatibility to overcome the loss of links.
In summary, this study makes two main contributions.
@@
  • We generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence.
@@
The paper is organized as follows. Section 2 presents the peak computation problem addressed in this paper and the main difficulty related to the search for an integer as an optimal solution of the problem. In Section 3, we introduce an illustrative example to which we apply the novel techniques in the following sections. This example is the simplest version of the Leslie model for studying the dynamics of a population decomposed into age classes. Then, in Section 4, we introduce the new class $\mathcal{KL}_{\rm gen}$. We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds on $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair. In Section 5, we present the concept of Opt-Lyapunov functions well-designed for our maximization problem. The classical comparison functions associated with classical Lyapunov functions are replaced by a new notion called compatibility. We also prove that the existence of a compatible Opt-Lyapunov function is equivalent to the existence of a useful pair. In Section 6, we summarize the results obtained, highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$, convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class, and Opt-Lyapunov functions.
@@ 2 Problem Statement And Motivated Example
2.1 The motivation : a discrete-time peak computation problem
2.1.1 Definition of the problem
@@
Let us consider a discrete-time dynamical system $(X^{\mathrm{in}},T)$ on $\mathbb{R}^{d}$ where
  • $X^{\mathrm{in}}$ is the nonempty subset of $\mathbb{R}^{d}$ of initial conditions;
  • $T$ is a nonzero self-map on $\mathbb{R}^{d}$ that updates the state-variable.
The system is autonomous and evolves without control, disturbances, or external inputs. In this paper, we are interested in the reachable values that maximize a given function. Then, given $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$, we define the associated discrete-time peak computation problem as follows:
$\begin{array}[]{lcl}\displaystyle{\operatorname*{Max}_{x,k}}&&\displaystyle{\varphi(T^{k}(x))}\\ &\text{ s. t.}&\displaystyle{x\in X^{\mathrm{in}},k\in\mathbb{N}}\end{array}$ (DPCP)
where $T^{k}$ is the $k$-fold composition of $T$ if $k$ is not null, and the identity if $k=0$.
Like any optimization problem, to solve a problem of the form (DPCP), we must compute
  • its optimal value, i.e., $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$;
  • its optimal solutions, i.e., couples $(\overline{x},\overline{k})\in X^{\mathrm{in}}\times\mathbb{N}$ such that
    $\varphi(T^{\overline{k}}(\overline{x}))=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}.$
An optimal solution $(x,k)$ of Problem (DPCP) is thus composed of an initial condition $x$ leading to a peak and its associated date $k$ of peak-realization.
@@
Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ for $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure. This is formulated in Assumption 1.
Assumption 1.
$\varphi(0)=0$.
2.1.2 A problem on integers
The optimal value of Problem (DPCP) can be described using a sequence of optimal values. Indeed, we can introduce a static classical optimization problem in which integer $k$ is fixed as follows:
$\begin{array}[]{lcl}\displaystyle{\operatorname*{Max}_{x}}&&\displaystyle{\varphi(T^{k}(x))}\\ &\text{ s. t.}&\displaystyle{x\in X^{\mathrm{in}}}\end{array}$ ($P_{k}$)
Then, we define the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ for all $k\in\mathbb{N}$ by:
$\nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}$ (1)
and its supremum:
$\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}.$ (2)
Naturally, if the optimal value of Problem  ($P_{k}$) is infinite, which is the same as $\nu_{k}=+\infty$ for some $k\in\mathbb{N}$, then the optimal value of Problem (DPCP) is infinite. We assume that the following holds: